OpenStudy (anonymous):

Use L'hopitals rule to find the limit. Limit as x approaches infinity. x sin (19/x)

OpenStudy (jamesj):

So to use l'Hopital's Rule, you will need to write your function as a ratio of two functions both of which have limit of 0 as x --> infty. Can you see how to write it that way?

OpenStudy (anonymous):

LIM X LIM SIN 19/X?

OpenStudy (jamesj):

\[ \lim_{x \to \infty} x = \infty \] not zero

OpenStudy (anonymous):

ok.

OpenStudy (anonymous):

so then the same with lim 19/infinity.

OpenStudy (jamesj):

To reiterate, we need to write \[ x \sin(19/x) = \frac{f(x)}{g(x)} \] where \[ \lim_{x \to \infty} f(x) = \lim_{x \to \infty} g(x) = 0 \]

OpenStudy (jamesj):

Then by l'Hopital's rule, provided the appropriate limits exist, \[ \lim_{x \to \infty} \frac{f(x)}{g(x)} = \lim_{x \to \infty} \frac{f'(x)}{g'(x)} \]

OpenStudy (jamesj):

So what f(x) and g(x) would do this?

OpenStudy (anonymous):

what would make g(x) and f(x) equal 0?

OpenStudy (jamesj):

not equal to zero, having limits equal to zero.

OpenStudy (jamesj):

hint: f(x) = sin(19/x). Now, what's g(x) ?

OpenStudy (jamesj):

If f(x) = sin(19/x) and f(x)/g(x) = x.sin(19/x), solve for g(x).

OpenStudy (anonymous):

1/sin (19/x)

OpenStudy (jamesj):

\[ \frac{f(x)}{g(x)} = \frac{\sin(19/x)}{g(x)} = x \sin(19/x) \] Therefore \( g(x) = 1/x \). Now, is it the case that the limits of f(x) and g(x) as x --> infty are both zero?

OpenStudy (anonymous):

they both get closer to 0

OpenStudy (anonymous):

they get smailler and smaller

OpenStudy (jamesj):

their limits are both zero. Now apply l'Hopital's rule. Calculate \[ \lim_{x \to \infty} \frac{f'(x)}{g'(x)} \]

OpenStudy (anonymous):

so the answer is 0

OpenStudy (jamesj):

No.

OpenStudy (anonymous):

f'(x) is -19 cos (19/x) / x^2

OpenStudy (jamesj):

Yes and g'(x) = ...?

OpenStudy (anonymous):

and the g'(x) is 1/x^2

OpenStudy (jamesj):

No, -1/x^2

OpenStudy (anonymous):

oops that what i meant.

OpenStudy (jamesj):

Hence f'(x)/g'(x) = ...?

OpenStudy (anonymous):

19 cos (19/x)

OpenStudy (jamesj):

Yes and therefore \[ \lim_{x \to \infty} \frac{f'(x)}{g'(x)} = ...? \]

OpenStudy (anonymous):

19

OpenStudy (jamesj):

Yes.

OpenStudy (anonymous):

sweet thanks sorry it took me a while. I really appreciate it!